OpenStudy (anonymous):

-x+y=1 x+y=11 solve please?

OpenStudy (anonymous):

Hmmmm.....same approach as I suggested for the previous problem. Here, adding will eliminate x and give an equation in y.

OpenStudy (anonymous):

can you walk me through it? i'm not so great at math

OpenStudy (anonymous):

-x + y = 1 x + y = 11 ---------- 2y = 12 y = 6 ---------- Using: x + y = 11 and y = 6, you get x + 6 = 11 x = 5

OpenStudy (anonymous):

hmm...i think i get it...

OpenStudy (anonymous):

grait explination